a mass of 88 grams of carbon dioxide will occupy a volume of how many liters at STP?
1
Expert's answer
2013-02-13T11:15:13-0500
STP is standard temperature and pressure, so you need to use PV = nRT
Pressure = 1 atm
Volume = unknown!
n = moles of CO2 = m/Mw = 88/44 = 2 moles
R = gas constant. 0.0821 L*atm / K*mol
Temperature = 273 K

V = nRT/P = 2*0.0821*273/1 = 44.82 L